# Astrometric constraints on stochastic gravitational wave background with neural networks

### v1

@@ Abstract
Astrometric measurements provide a unique avenue for constraining the stochastic gravitational wave background (SGWB). In this work, we investigate the application of two neural network architectures, a fully connected network and a graph neural network, for analyzing astrometric data to detect the SGWB. Specifically, we generate mock Gaia astrometric measurements of the proper motions of sources and train two networks to predict the energy density of the SGWB, $\Omega_{\text{GW}}$. We evaluate the performance of both models under varying input datasets to assess their robustness across different configurations. Our results demonstrate that neural networks can effectively measure the SGWB, showing promise as tools for addressing systematic uncertainties and modeling limitations that pose challenges for traditional likelihood-based methods.
@@ I Introduction
Cosmological sources of the SGWB include primordial processes such as inflation [1], cosmic strings [2], phase transitions [3] in the early universe. In contrast, astrophysical sources contribute through events like compact binary mergers, supernovae, and rotating neutron stars [4]. The study of the SGWB is crucial, as it provides unique insights into the early universe, fundamental physics, and astrophysical populations and processes. Detecting and characterizing the SGWB could validate inflationary models, probe the physics of the early universe, and shed light on compact object populations across cosmic time [5, 6, 7].
Determining the origin of the SGWB is a challenging task, requiring its characterization across a broad frequency range. Given the multitude of potential contributing sources, it is crucial to probe the SGWB amplitude at various frequencies. At high frequencies, future technological advancements and missions, including space-based observatories [8, 9], ground-based detector networks [10], and ultra-high-frequency GW experiments [11], will greatly enhance sensitivity. In contrast, the low-frequency regime will be investigated using cosmological and astrophysical observations, such as the Cosmic Microwave Background (CMB) B-mode polarization measurements [12] and Pulsar Timing Arrays (PTAs) [13], with astrometry also playing a pivotal role.
@@
However, note that currently the PTA measurements [13] (around $10^{-9}$ Hz), the joint CMB+BBN constraint on the relativistic energy density [25] (for frequencies higher than $10^{-10}$ Hz), and the CMB $\mu$-distortion constraints from COBE/FIRAS [26] (between $10^{-16}$ and $10^{-9}$ Hz) provide stronger limits on the SGWB amplitude. Nevertheless, with the anticipated high-precision proper motion measurements from the upcoming series of Gaia data release, as well as the Nancy Grace Roman Space Telescope [27, 28] and proposed upgrade mission THEIA [29, 30], astrometry has the potential to become a competitive tool for constraining the SGWB within this frequency window.
@@
In this work, we explore the potential of neural networks (NNs) to analyze astrometric data and constrain the energy density of the SGWB. Traditional likelihood-based methods face several challenges in the analysis and do not fully leverage Gaia’s astrometric measurements. First, due to the complexities in modeling the intrinsic proper motion of stars in our galaxy, current analyses are limited to distant quasars. While quasars offer the advantage of stability in their motion, their numbers are significantly smaller than the stars within our galaxy. Moreover, even with stable quasar catalogs, systematic errors in astrometric measurements and source misidentification remain major obstacles [23].
@@
To fully realize the potential of astrometric surveys, it is crucial to increase the number of available sources by incorporating measurements of all stars in our galaxy. Achieving this goal entails overcoming several challenges: accurately modeling galactic rotation to subtract stars’ intrinsic proper motions, establishing robust sample selection criteria, and addressing the substantial computational demands associated with analyzing billions of stars. Finally, future data releases will provide time-series measurements, offering significant advantages for constraining GWs [31]. However, this will also substantially increase the complexity of the analysis. Addressing these issues will be critical to fully utilizing the power of astrometric data.
@@
The application of NNs to cosmological observations has rapidly developed across various experimental datasets, demonstrating significant advantages and driving revolutions in data analysis in the era of big data. The flexibility of NNs has the potential to overcome the difficulties and limitations we face with traditional methods. Our ultimate aim is to develop architectures capable of accurately measuring the SGWB using the vast amount of future astrometric data. As a first step, for the first time, we test the application of NNs to GW astrometry by using simulated Gaia quasar mock catalogs. In this preliminary study, we aim to lay the groundwork for future extensions and applications to real data.
@@
In this work, we design two types of NNs that take astrometric data (positions, proper motions, etc.) as input and output a prediction of the SGWB amplitude, $\Omega_{\text{GW}}$. The first type is a Fully Connected Network (FCN), known for its simplicity and flexibility. The second is a Graph Neural Network (GNN), designed to process graph-structured data and capture relationships between elements. We place particular emphasis on the GNN due to the natural alignment of our dataset with a graph-based representation.
@@
This paper is organized as follows. In Sec. II, we provide an overview of the theoretical framework underlying GW astrometry and describe the details of simulating mock datasets. In Sec. III, we outline the architectures of the FCN and GNN used in this study. In Sec. IV, we present the results of testing the performance of both architectures. First, we evaluate their performance using homogeneously distributed sources, followed by tests in inhomogeneous cases with galactic masks applied. Finally, we conclude in Sec. V.
@@ II Theoretical formalism and creation of mock data
where $\rho_{\rm GW}$ is energy density of GWs and $\rho_{c}$ is the critical density for a flat Universe,
@@
In this paper, we want to address constraints on the SGWB coming from astrometric measurements, due to the deflection caused by GWs along the light trajectories of sources in the sky. It has been shown that the expected upper bound on the energy density $\Omega_{\text{GW}}$ is
@@
where $\Delta\mu^{2}$ is the variance associated to the proper motion at a given frequency, and $N$ the total number of sources. The magnitude of the proper motion $\mu$ is often decomposed as
@@
To train the NNs and evaluate their performance through testing and validation, we create quasar mock catalogs using the Python package provided by the Gaia collaboration, pygaia [34]. Quasars are a suitable choice for the first step, as they have negligible intrinsic proper motion compared to other types of sources. Assuming a homogeneous distribution of sources, we randomly generate the right ascension and declination coordinates. The errors in the astrometric measurements depend on the brightness of the sources. Therefore, we randomly assign a G-band magnitude value for each source, drawn from a uniform distribution between $20.716<G<$ 20.7. The upper value corresponds to the nominal magnitude limit of Gaia, while the lower bound is determined by the quasar catalog presented in [35]. Then the proper_motion_uncertainty function from pygaia.errors.astrometric provides the associated errors for both the declination and right ascension components of each source, based on a given magnitude and the assumed Gaia data release. We assume the Gaia DR3 sensitivity. The details of the modeling for astrometric uncertainties can be found in the documentation provided by Gaia [36].
We then inject the SGWB signal, specifically the quadrupole component of the vector harmonics [37]. To achieve this, we randomly generate the values of the multipole coefficients of the vector harmonics (see [37, 22] for the full expressions) from a uniform distribution over $[0,1]$ and renormalize the amplitude using the relation Eq. (3), which connects the quadrupole power $P_{2}$ to the SGWB amplitude $\Omega_{\text{GW}}$ as
@@
For training the NNs, the input set of parameters is characterized by six features: right ascension and declination coordinates, proper motion components for right ascension and declination, and the error amplitudes for right ascension and declination estimated from the magnitudes of the sources. We generate a set of mock catalogs, each with a different realization of noise and varying amplitudes of the SGWB injection, which are randomly selected from the linear uniform distribution $[0,1]$. This range is made to reflect the typical sensitivity of the Gaia DR3 data, which spans between $\Omega_{\text{GW}}\sim 0.01$ to $0.1$ depending on the number of sources. The NNs are trained to predict the correct values of $\Omega_{\text{GW}}$ based on the input astrometric measurements. We test five cases with different numbers of sources, chosen as $N_{s}=500,1000,2000,6000$ and $12000$. For each configuration with a fixed value of $N_{s}$, we generate 8000 mock simulated datasets and train the NNs independently for each configuration.
To demonstrate an example where NNs can showcase their flexibility, we consider an inhomogeneous distribution of sources. One potential cause of inhomogeneities in quasar samples is the masking of the galactic plane. To simulate such a case, we exclude sources in a symmetric central band around the celestial equator. Specifically, the declination coordinates of the sources are randomly and uniformly distributed above and below this equatorial exclusion zone, defined by the range $[-\delta,+\delta]$. We test four cases with $\delta=10^{\circ}$, $20^{\circ}$, $40^{\circ}$, and $60^{\circ}$.
@@ III Neural network architectures
The application of machine learning techniques to the analysis and interpretation of GW data represents a significant advancement in the field. These algorithms are particularly well-suited for tackling complex problems, such as pattern recognition in large and intricate datasets, feature extraction, and significantly improving computation time.
@@
The second type of NN we focus on in this work is GNN [38, 39], a deep NN developed to handle graph-structured data. A graph is a structure made by nodes and relationships between nodes represented by connections, called edges, while information is encoded in features associated with nodes or edges and used for predictions. Nowadays, graph-structured data can be found extensively among various scientific and social disciplines, such as chemistry, social networks, cosmology, making it important to deep learning models to handle such data. Consequently, interest in graph representation learning has grown significantly in recent years.
In this work, the NN architectures are implemented by using PyTorch framework. The training and validation datasets, consisting of both the mock data and their corresponding target values, are prepared by splitting the original dataset into 80% for training and 20% for validation.
@@
Target labels and mock data are firstly normalized, which helps to improve the stability and efficiency of the NNs training process. The implemented FCN architecture consists of three linear layers, where the first two are followed by a ReLU activation function [40]. The input size is the number of features (six in our case), while each hidden layer contains 64 neurons. The final output returns a single value through the last linear layer, which represents the prediction of the network for a regression task. Moreover, to prevent overfitting, a dropout layer with probability of 0.5 is added at the end of the forward process.
The loss function used is the MSELoss, that measures the mean squared error between each element in the prediction and true values, while Adam optimizer [41] is used, with learning rate of $10^{-4}$. To regularize the training, we implement an early stopping technique, to stop the training if the validation loss does not improve by a certain tolerance over 100 epochs.
Figure 1: Illustrative graph representation for the different number of sources. The number of edges for each graph has been chosen to ensure that almost all points are connected. In this plot, the corresponding number of edges for each configuration is $n_{\text{ed},500}=3270$, $n_{\text{ed},1000}=6212$, $n_{\text{ed},2000}=12493$, $n_{\text{ed},6000}=36976$, $n_{\text{ed},12000}=75944$, respectively.
@@
Among all the different types of GNN tasks, graph classification is used to classify entire graphs into various categories, based on structural graph features of a given dataset of graphs. We use PyTorch Geometric [42].
Firstly, we create graphs as input for the GNN. In our graph structure each node is represented by a source position, with six features. We evaluate the Haversine distance between each pair of points and an edge between two nodes is formed if their distance is less than a given threshold. To ensure that each node is connected to its neighbors, the threshold is chosen slightly larger than the average distance. The average distance scales as $\propto N_{s}^{-1/2}$ and the distance threshold we used are $0.317$ rad, $0.224$ rad, $0.159$ rad, $0.091$ rad and $0.065$ rad for $N_{s}=500$, $N_{s}=1000$, $N_{s}=2000$, $N_{s}=6000$, and $N_{s}=12000$, respectively. In this way, we can preserve the local structural relationships necessary for accurate predictions without introducing unnecessary complexity.
To store the graph data, we use Data tool from PyTorch Geometric. It takes node features, graph connectivity, positions and labels (in this case one value of $\Omega_{\text{GW}}$ for the entire graph), and constructs the graph representation for a given dataset. The chosen numbers of sources and realizations are the same of the FCN. To illustrate what a single representation looks like, we plot the graph representation in three dimension for different number of sources in Fig. 1.
The architecture consists of five convolution layers, by using GCNConv [43] implemented in PyTorch models, followed by a final linear layer to produce a single output value for each graph. ReLU activations are applied after each graph convolution to introduce non-linearity. At the end of the forward process, a global mean pooling operation combines node embeddings into a graph-level representation, using a global mean pooling operation. Finally, a dropout layer with probability of 0.5 is added, and the linear layer outputs the final prediction. The input size corresponds to the number of features, while each hidden layer contains 128 neurons.
@@
Figure 3: Comparison of GNN results for different numbers of layers. Predictions are produced in the same way as in Fig. 2, where the GNN has five layers. The left panel shows the results from a GNN with only two convolutional layers, while the right panel shows the results from a GNN with three convolutional layers in its architecture.
The training process uses Adam optimizer with learning rate of $10^{-4}$ and MSE loss function. We implement an early stopping technique, with patience of 100 epochs.

### v2

@@ Abstract
Astrometric measurements provide a unique avenue for constraining the stochastic gravitational wave background (SGWB). In this work, we investigate the application of two neural network architectures, a fully connected network and a graph neural network, for analyzing astrometric data to detect the SGWB. Specifically, we generate mock Gaia astrometric measurements of the proper motions of sources and train two networks to predict the energy density of the SGWB, $\Omega_{\text{GW}}$. We evaluate the performance of both models under varying input datasets to assess their robustness across different configurations. We also perform a direct comparison with a likelihood-based approach using Markov chain Monte Carlo (MCMC) methods, finding out that the neural-network-based approach is significantly faster, taking on the order of minutes, compared to MCMC’s order of days, while still capturing the same features in the data. Our results demonstrate that neural networks can effectively constrain the SGWB, showing promise as tools for addressing systematic uncertainties and modeling limitations that pose challenges for traditional likelihood-based methods.
@@ I Introduction
Cosmological sources of the SGWB include primordial processes such as inflation [1], cosmic strings [2], and phase transitions [3] in the early Universe. In contrast, astrophysical sources contribute through events like compact binary mergers, supernovae, and rotating neutron stars [4]. The study of the SGWB is crucial, as it provides unique insights into the early Universe, fundamental physics, and astrophysical populations and processes. Detecting and characterizing the SGWB could validate inflationary models, probe the physics of the early Universe, and shed light on compact object populations across cosmic time [5, 6, 7].
Determining the origin of the SGWB is a challenging task requiring its characterization across a broad frequency range. Given the multitude of potential contributing sources, it is crucial to probe the SGWB amplitude at various frequencies. Future technological advancements and missions will probe a wide range of frequencies ranging from $10^{-3}$ to $10^{8}$Hz, from space-based observatories [8, 9] and ground-based detector networks [10] to ultrahigh-frequency GW experiments [11], substantially improving sensitivity. In contrast, frequencies lower than $10^{-8}$Hz can be investigated using cosmological and astrophysical observations, such as the cosmic microwave background (CMB) B-mode polarization measurements [12] and pulsar timing arrays (PTAs) [13], with astrometry also playing a pivotal role.
@@
However, note that currently the PTA measurements [13] (around $10^{-9}$ Hz), the joint CMB+BBN constraint on the relativistic energy density [25] (for frequencies higher than $10^{-10}$ Hz), and the CMB $\mu$-distortion constraints from COBE/FIRAS [26] (between $10^{-16}$ and $10^{-9}$ Hz) provide stronger limits on the SGWB amplitude. Nevertheless, with the anticipated high-precision proper motion measurements from the upcoming series of the Gaia data release, as well as the Nancy Grace Roman Space Telescope [27, 28] and proposed upgrade mission THEIA [29, 30], astrometry has the potential to become a competitive tool for constraining the SGWB within this frequency window.
@@
In this work, we explore the potential of neural networks (NNs) to analyze astrometric data and constrain the energy density of the SGWB. Traditional likelihood-based methods face several challenges in the analysis and do not fully leverage Gaia’s astrometric measurements. First, due to the complexities in modeling the intrinsic proper motion of stars in our galaxy, current analyses are typically limited to distant quasars. While quasars offer the advantage of stability in their motion, their numbers are significantly smaller than the stars within our galaxy. Moreover, even with stable quasar catalogs, systematic errors in astrometric measurements and source misidentification remain major obstacles [23].
@@
To fully realize the potential of astrometric surveys, it is crucial to increase the number of available sources by incorporating measurements of all stars in our galaxy. Achieving this goal entails overcoming several challenges: accurately modeling galactic rotation to subtract stars’ intrinsic proper motions, establishing robust sample selection criteria, and addressing the substantial computational demands associated with analyzing billions of stars. Finally, future data releases will provide time-series measurements, offering significant advantages for constraining GWs [31]. Additionally, synergy with PTA data presents an interesting avenue for further exploration [32, 33, 34, 35]. However, this will also substantially increase the complexity of the analysis. Addressing these issues will be critical to fully utilizing the power of astrometric data.
@@
The application of NNs to cosmological observations has rapidly developed across various experimental datasets, demonstrating significant advantages and driving revolutions in data analysis in the era of big data. The flexibility of NNs has the potential to overcome the difficulties and limitations we face with traditional methods. Our ultimate aim is to develop architectures capable of accurately constraining the SGWB using the vast amount of future astrometric data. As a first step, for the first time, we test the application of NNs to GW astrometry by using simulated Gaia quasar mock catalogs. In this preliminary study, we aim to lay the groundwork for future extensions and applications to real data.
@@
In this work, we design two types of NNs that take astrometric data (positions, proper motions, etc.) as input and output a prediction of the SGWB amplitude, $\Omega_{\text{GW}}$. The first type is a fully connected network (FCN) known for its simplicity and flexibility. The second is a graph neural network (GNN) designed to process graph-structured data and capture relationships between elements. We place particular emphasis on the GNN due to the natural alignment of our dataset with a graph-based representation.
@@
This paper is organized as follows. In Sec. II, we provide an overview of the theoretical framework underlying GW astrometry and describe the details of simulating mock datasets. In Sec. III, we outline the architectures of the FCN and GNN used in this study. In Sec. IV, we present the results of testing the performance of both architectures. First, we evaluate their performance using homogeneously distributed sources, followed by tests in inhomogeneous cases with galactic masks applied. Moreover, we perform a direct comparison with the standard Markov chain Monte Carlo (MCMC) methods. Finally, we conclude in Sec. V.
@@ II Theoretical formalism and creation of mock data
where $\rho_{\rm GW}$ is the energy density of GWs and $\rho_{c}$ is the critical density for a flat universe,
@@
In this paper, we want to address constraints on the SGWB coming from astrometric measurements, due to the deflection caused by GWs along the light trajectories of sources in the sky. It has been shown that the expected upper bound on the energy density $\Omega_{\text{GW}}$ is [16]
@@
where $\Delta\mu^{2}$ is the variance associated with the proper motion at a given frequency, and $N$ the total number of sources. The magnitude of the proper motion $\mu$ is often decomposed as
@@
To train the NNs and evaluate their performance through testing and validation, we create quasar mock catalogs using the Python package provided by the Gaia Collaboration, pygaia [38]. Quasars are a suitable choice for the first step, as they have negligible intrinsic proper motion compared to other types of sources. Assuming a homogeneous distribution of sources, we randomly generate the right ascension and declination coordinates. The errors in the astrometric measurements depend on the brightness of the sources. Therefore, we randomly assign a G-band magnitude value [39, 40] for each source, drawn from a uniform distribution between $16<G<20.7$. The upper value corresponds to the nominal magnitude limit of Gaia, while the lower bound is determined by the quasar catalog presented in [39]. Then the proper_motion_uncertainty function from pygaia.errors.astrometric provides the associated errors for both the declination and right ascension components of each source, based on a given magnitude and the assumed Gaia data release. We assume the Gaia DR5 sensitivity. The details of the modeling for astrometric uncertainties can be found in the documentation provided by Gaia [40].
We then inject the SGWB signal, specifically the quadrupole component of the vector harmonics [41]. To achieve this, we randomly generate the values of the multipole coefficients of the vector spherical harmonics (see [41, 22] for the full expressions) from a uniform distribution over $[0,1]$ and renormalize the amplitude using the relation Eq. (3), which connects the quadrupole power $P_{2}$ to the SGWB amplitude $\Omega_{\text{GW}}$ as
@@
For training the NNs, the input set of parameters is characterized by six features: right ascension and declination coordinates, proper motion components for right ascension and declination, and the error amplitudes for right ascension and declination estimated from the magnitudes of the sources. We generate a set of mock catalogs, each with a different realization of noise and varying amplitudes of the SGWB injection, which are randomly selected from the linear-uniform distribution $[0,1]$. The NNs are trained to predict the correct values of $\Omega_{\text{GW}}$ injected based on the input astrometric measurements. We test three cases with different numbers of sources, chosen as $N_{s}=500,\,2000,\,12000$. For each configuration with a fixed value of $N_{s}$, we generate 8000 mock simulated datasets and train the NNs independently for each configuration.
To demonstrate an example where NNs can showcase their flexibility, we consider an inhomogeneous distribution of sources. One potential cause of inhomogeneities in quasar samples is the masking of the galactic plane. To simulate such a case, we exclude sources in a symmetric central band around the celestial equator. Specifically, the declination coordinates of the sources are randomly and uniformly distributed above and below this equatorial exclusion zone, defined by the range $\delta,+\delta]$. We test four cases with $\delta=10^{\circ}$, $20^{\circ}$, $40^{\circ}$, and $60^{\circ}$.
@@ III Neural network architectures
The application of machine learning techniques to the analysis and interpretation of GW data represents a significant advancement in the field. These algorithms are particularly well suited for tackling complex problems, such as pattern recognition in large and intricate datasets, feature extraction, and significantly improving computation time.
@@
The second type of NN we focus on in this work is GNN [42, 43], a deep NN developed to handle graph-structured data. A graph is a structure made by nodes and relationships between nodes represented by connections called edges, while information is encoded in features associated with nodes or edges and used for predictions. Nowadays, graph-structured data can be found extensively among various scientific and social disciplines, such as chemistry, social networks, and cosmology, making it important to deep learning models to handle such data. Consequently, interest in graph representation learning has grown significantly in recent years.
In this work, the NN architectures are implemented via the PyTorch framework. Specifically, we have carried out several tests to determine the best-performing architectures and hyperparameter settings for the networks, leveraging Optuna for optimization [44]. The training and validation datasets consisting of both the mock data and their corresponding target values are prepared by splitting the original dataset into 80% for training and 20% for validation [45].
Figure 1: Illustrative graph representation for the different number of sources. The number of edges for each graph has been chosen to ensure that almost all points are connected. In this plot, the corresponding number of edges for each configuration is $n_{\text{ed},500}=3062$, $n_{\text{ed},2000}=12758$, $n_{\text{ed},12000}=75746$, respectively.
@@
The input mock data are normalized by subtracting the mean and dividing by the standard deviation computed across the samples and their features, which helps to improve the stability and efficiency of the NNs training process [46]. The implemented FCN architecture consists of five hidden linear layers, each followed by a ReLU activation function [47]. The input size corresponds to the number of features (six in our case), and each hidden layer contains 256 neurons. After the final hidden layer, the output is averaged along the feature dimension to aggregate the learned representations. To reduce overfitting, dropout with a probability of 0.4 is applied. The resulting vector is then passed through a final linear layer that outputs a single scalar value, representing the network prediction for a regression task.
The loss function used is the MSELoss that measures the mean squared error (MSE) between each element in the prediction and true values, while the Adam optimizer [48] is used with a learning rate of $10^{-3}$. To regularize the training, we implement an early stopping technique to stop the training if the validation loss does not improve by a certain tolerance over 100 epochs.
@@
Among all the different types of GNN tasks, graph classification is used to classify entire graphs into various categories, based on structural graph features of a given dataset of graphs.
First, we create a set of graphs as input for the GNN. Each realization is converted into a graph, where nodes correspond to sources characterized by six astrophysical features. The normalization is the same as for the FCN training dataset. Edges between nodes are defined based on angular proximity. Specifically, we compute the Haversine distance between pairs of source positions and form an undirected edge if the distance falls below a chosen threshold. To ensure that each node is connected to its local neighborhood, the threshold is chosen slightly larger than the average distance, which scales as $\propto N_{s}^{-1/2}$, where $N_{s}$ is the number of sources. In our implementation, we use thresholds of $0.317$, $0.159$, and $0.065$ rad for $N_{s}=500$, $N_{s}=2000$, and $N_{s}=12000$, respectively. This procedure preserves the relevant local structure while avoiding unnecessary complexity.
To store the graph data, we use Data tool from PyTorch Geometric [49], which contains the node features, edge connectivity, and a single graph-level label corresponding to the simulated value of $\Omega_{\text{GW}}$. The numbers of sources and realizations match those of the FCN. A three-dimensional visualization of example graphs for different $N_{s}$ values is shown in Fig. 1.
The architecture consists of five graph convolutional layers implemented with GCNConv [50]. Each convolutional layer is followed by a ReLU activation to introduce non-linearity. The input size corresponds to the number of node features, while each hidden layer contains 128 channels. After the final convolution, node embeddings are aggregated into graph-level representations using a global mean pooling operation, which averages over all nodes in a graph. To mitigate overfitting, a dropout layer with a probability of 0.4 is applied to the pooled representations. A final linear layer maps this representation to a single scalar output, yielding the predicted value of $\Omega_{\mathrm{GW}}$.
@@
Training is performed using the Adam optimizer with a learning rate of $10^{-4}$ and MSE loss. Early stopping with a patience of 100 epochs and tolerance of $10^{-4}$ is employed to prevent overfitting.